OpenStudy (anonymous):

If a line has a slope of 2 and passes though (4,-1) what is the y intersept

OpenStudy (anonymous):

Plug in your values. y=mx+b (m = 2, x = 4, y = -1) y=mx+b -1=2(4)+b -1=8+b b=-9 The y intercept is (0, -9)